OpenStudy (sleepyjess):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 3 + 3 + 9 + ... + 81

OpenStudy (sleepyjess):

@e.mccormick

OpenStudy (sleepyjess):

darn, clicked the wrong thing...

OpenStudy (freckles):

closed means you succeeded

OpenStudy (freckles):

lol

OpenStudy (freckles):

oh so no

OpenStudy (freckles):

so hmm we need to find a pattern

OpenStudy (sleepyjess):

I meant to click delete, but instead clicked close because it moved right as I clicked lol

OpenStudy (freckles):

so it really goes like this...-9,-3,3,9,...,81

OpenStudy (freckles):

it looks like this is an arithmetic sequence

OpenStudy (freckles):

do you see the common difference?

OpenStudy (freckles):

-3-(-9)=? 3-(-3)=? 9-3=? all of these are equal to what?

OpenStudy (sleepyjess):

adding 6

OpenStudy (freckles):

yep do first I'm going to play with this sequence then I will go back to the series you mentioned

OpenStudy (freckles):

\[a_n=a_1+d(n-1)\] d is the common difference and a_1 is the first term in the sequence

OpenStudy (freckles):

you already identified d

OpenStudy (freckles):

as 6

OpenStudy (freckles):

now a_1 is ?

OpenStudy (sleepyjess):

-9 so \(a_n\)=-9+6(n-1)

OpenStudy (freckles):

right we need to figure out for what number n is the last term 81

OpenStudy (freckles):

\[81=-9+6(n-1)\] a_n is 81 but what is n well solve this for n to find out :)

OpenStudy (sleepyjess):

81 = -9 + 6n - 6 81 = -15 + 6n 96 = 6n 16 = n

OpenStudy (freckles):

\[\text{ your answer is going \to be \in this form } \sum_{n=1}^{16}a_n\] last step just replace a_n with what you found above

OpenStudy (sleepyjess):

The -9 + 6n?

OpenStudy (freckles):

-9+6(n-1)

OpenStudy (sleepyjess):

hmmm... I guess I will go with the option closest to that... which isn't really any of them o.O

OpenStudy (freckles):

\[\sum_{n=1}^{16}a_n=\sum_{n=1}^{16}(-9+6(n-1))\] you could combine like terms and there are also other ways to write the summation but you are done you did what it said which was write as a summutation

OpenStudy (freckles):

Well let me ask you a question

OpenStudy (freckles):

do they start it at n=0 ...

OpenStudy (freckles):

or some other number

OpenStudy (sleepyjess):

I don't know what they start it at, that was all the information given :/

OpenStudy (freckles):

oh i thought they gave you choices

OpenStudy (freckles):

it sounded like you were saying oyu have to select an option

OpenStudy (sleepyjess):

oh, yes, they give choices http://prntscr.com/68pk8z

OpenStudy (freckles):

so all the choices start off with n=0

OpenStudy (freckles):

do you see that

OpenStudy (freckles):

ours starts off at n=1

OpenStudy (freckles):

we can manipulate that

OpenStudy (freckles):

\[\sum_{n=1}^{16}a_n=\sum_{n=1}^{16}(-9+6(n-1)) \\ \\ \sum_{n=1}^{n=16}(-9+6(n-1)) \\ \sum_{n-1=0}^{n-1=15}(-9+6(n-1))\]

OpenStudy (freckles):

now replace n-1 wherever you see n-1 with new n

OpenStudy (freckles):

\[\sum_{n=0}^{n=15}(-9+6(n))\]

OpenStudy (sleepyjess):

that makes more sense :) thanks for coming to help even though I accidentally closed it haha

OpenStudy (freckles):

do you see how i manipulated like we had from n=1 to n=16 I subtracted 1 on both sides for these equations n=1 n-1=1-1 n-1=0 and n=16 n-1=16-1 n-1=15 and our expression was already in terms of (n-1)

OpenStudy (sleepyjess):

yes

OpenStudy (freckles):

cool stuff

OpenStudy (freckles):

also the bottom two make no since because we do not have an infinite series :) and the second one also doesn't make sense because none of the beginning terms in your sequence are multiples of -54 so I think you could have done less work here and just used process of elimination a long with common sense but it is also nice to see how to set it up :)

OpenStudy (sleepyjess):

I figured it was the first one, but I didn't want to just pick an option, I wanted to know how to do it

OpenStudy (freckles):

And I totally think that is cool :)